OpenStudy (anonymous):

Solve the system algebraically. 3x - 2y - 1 = 0 y = 5x + 4 What is the solution? {(-9/7,-17/7)} {(9/7,17/7)} {(9/7,-17/7)}

OpenStudy (anonymous):

3x-2(5x+4)-1=0=>-7x=9=>x=-9/7 y=5(-9/7)+4=-17/7

OpenStudy (anonymous):

top one

OpenStudy (anonymous):

?

OpenStudy (lalaly):

yes